OpenStudy (anonymous):

One circular table has a diameter of 9 ft, and another circular table has a diameter of 20 ft. How much greater is the area of the larger table? Round to the nearest whole number

OpenStudy (anonymous):

91 * 3.14 = 274

OpenStudy (anonymous):

about 1,002 ft2 about 251 ft2 about 146 ft2 about 336 ft2

OpenStudy (anonymous):

which one

OpenStudy (anonymous):

probably 251

OpenStudy (anonymous):

lets double check

OpenStudy (anonymous):

Area = pi * r^2

OpenStudy (anonymous):

r is radius. diameter is twice the size of the radius. So a circle with diameter of 20 will have a radius of 10 and a circle with a diameter of 9 will have a radius of 4.5.

OpenStudy (anonymous):

to find the answer subtract the areas of the two circles

OpenStudy (anonymous):

pi * 10^2 - pi * 4.5^2

OpenStudy (anonymous):

pi * 100 - pi * 20.25

OpenStudy (anonymous):

pi * 79.75 = 3.14 * 79.75 = 250.415

OpenStudy (anonymous):

so the answer is closest to 251 ft^2